# Sharpness of Minima in Deep Matrix Factorization: Exact Expressions

### v4

@@ D.2 Experiment Details and Additional Experiments
(a) Trajectories of GD initialized at $\bm{w}_{0}$ with step sizes $\geq 2/{\lambda_{\max}}$ are depicted by colored lines, and their corresponding convergence points are marked by colored $\star$ symbols.
(b) Trajectories of GD initialized at $\bm{w}_{0}$ with step sizes $2/{\lambda_{\max}}$ are depicted by colored lines, and their corresponding convergence points are marked by colored $\star$ symbols.
Figure 3: Contour map of the loss landscape around a minimum in a 15-layer overparameterized scalar factorization of a random scalar. GD with different step sizes $\eta$, indicated by different colors, is initialized within a radius of $10^{-9}$ from the minimum, in the direction of the Hessian eigenvector corresponding to the largest eigenvalue. The vector $\bm{v}_{1}$ denotes the eigenvector of the Hessian corresponding to the largest eigenvalue, while $\bm{v}_{N}$ denotes the eigenvector corresponding to the smallest eigenvalue. The value of $\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{w}^{*}))$ is computed using the closed-form expression derived in Theorem 21.
For the experiment, we first generate the layer dimensions randomly and then construct the optimal layers $\bm{W}_{1}^{*},\bm{W}_{2}^{*},\dots,\bm{W}_{L}^{*}]$ as Gaussian random matrices, with each entry sampled from $N(0,1)$ according to the generated dimensions. Then, we compute $\bm{M}$ or $m$ by $\prod_{j=1}^{L}\bm{W}^{*}_{j}$. We then perform the same experiments as in Figs. 12, varying the depth, dimensions, and initialization distance $r$ (as shown in Figs. 3-7). We note that oscillations occur along the eigenvector corresponding to the maximum eigenvalue of the Hessian. The dimensions of the factors, i.e., $d_{0},d_{1},\dots,d_{L}$}, in Fig. 6 are given by 1,9,4,8,24,16,17,11,21,3,22,3,3,15,3,18,17,16,5,12,1{1,9,4,8,24,16,17,11,21,3,22,3,3,15,3,18,17,16,5,12,1}, which implies $N=2421$, while the dimensions of the factors in Fig. 7 are given by 1,9,4,8,1{1,9,4,8,1}, which implies $N=293$.
(a) Normalized training error across iterations, i.e., $\mathcal{L}(\bm{w}_{k})/\lVert\bm{M}\rVert_{F}^{2}$.
(b) Normalized $\ell^{2}$ distance of $\bm{w}_{k}$ from the minimum $\bm{w}^{*}$, i.e, $\norm{\bm{w}_{k}-\bm{w}^{*}}_{2}^{2}/\norm{\bm{w}^{*}}_{2}^{2}$.
@@
(d) Normalized training error across iterations, i.e., $\mathcal{L}(\bm{w}_{k})/\lVert\bm{M}\rVert_{F}^{2}$.
(e) Normalized $\ell^{2}$ distance of $\bm{w}_{k}$ from the minimum $\bm{w}^{*}$, i.e, $\norm{\bm{w}_{k}-\bm{w}^{*}}_{2}^{2}/\norm{\bm{w}^{*}}_{2}^{2}$.
@@
Figure 4: GD dynamics with different step sizes indicated by different colors for general matrix factorization, $\bm{M}=\bm{L}\bm{R}^{\top}$, of a random Gaussian matrix, where $\bm{L}\in\mathbb{R}^{10\times 30}$ and $\bm{R}\in\mathbb{R}^{20\times 30}$. The value of $\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{w}^{*}))$ is computed using the closed-form expression derived in Corollary 25.
(a) Normalized training error across iterations, i.e., $\mathcal{L}(\bm{w}_{k})/\lVert\bm{M}\rVert_{F}^{2}$.
(b) Normalized $\ell^{2}$ distance of $\bm{w}_{k}$ from the minimum $\bm{w}^{*}$, i.e, $\norm{\bm{w}_{k}-\bm{w}^{*}}_{2}^{2}/\norm{\bm{w}^{*}}_{2}^{2}$.
@@
(d) Normalized training error across iterations, i.e., $\mathcal{L}(\bm{w}_{k})/\lVert\bm{M}\rVert_{F}^{2}$.
(e) Normalized $\ell^{2}$ distance of $\bm{w}_{k}$ from the minimum $\bm{w}^{*}$, i.e, $\norm{\bm{w}_{k}-\bm{w}^{*}}_{2}^{2}/\norm{\bm{w}^{*}}_{2}^{2}$.
@@
Figure 5: GD dynamics with different step sizes indicated by different colors for general matrix factorization, $\bm{M}=\bm{L}\bm{R}^{\top}$, of a random Gaussian matrix, where $\bm{L}\in\mathbb{R}^{25\times 30}$ and $\bm{R}\in\mathbb{R}^{20\times 30}$. The value of $\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{w}^{*}))$ is computed using the closed-form expression derived in Corollary 25.
(a) Normalized training error across iterations, i.e., $\mathcal{L}(\bm{w}_{k})/m^{2}$.
(b) Normalized $\ell^{2}$ distance of $\bm{w}_{k}$ from the minimum $\bm{w}^{*}$, i.e, $\norm{\bm{w}_{k}-\bm{w}^{*}}_{2}^{2}/\norm{\bm{w}^{*}}_{2}^{2}$.
@@
(d) Normalized training error across iterations, i.e., $\mathcal{L}(\bm{w}_{k})/m^{2}$.
(e) Normalized $\ell^{2}$ distance of $\bm{w}_{k}$ from the minimum $\bm{w}^{*}$, i.e, $\norm{\bm{w}_{k}-\bm{w}^{*}}_{2}^{2}/\norm{\bm{w}^{*}}_{2}^{2}$.
@@
Figure 6: GD dynamics with different step sizes indicated by different colors are initialized within a radius of $10^{-9}$ from the minimum in the direction of the Hessian eigenvector corresponding to the largest eigenvalue, for a 20-layer overparameterized scalar factorization of a random scalar. The value of $\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{w}^{*}))$ is computed using the closed-form expression derived in Theorem 21.
(a) Normalized training error across iterations, i.e., $\mathcal{L}(\bm{w}_{k})/m^{2}$.
(b) Normalized $\ell^{2}$ distance of $\bm{w}_{k}$ from the minimum $\bm{w}^{*}$, i.e, $\norm{\bm{w}_{k}-\bm{w}^{*}}_{2}^{2}/\norm{\bm{w}^{*}}_{2}^{2}$.
@@
(d) Normalized training error across iterations, i.e., $\mathcal{L}(\bm{w}_{k})/m^{2}$.
(e) Normalized $\ell^{2}$ distance of $\bm{w}_{k}$ from the minimum $\bm{w}^{*}$, i.e, $\norm{\bm{w}_{k}-\bm{w}^{*}}_{2}^{2}/\norm{\bm{w}^{*}}_{2}^{2}$.
@@
Figure 7: GD dynamics with different step sizes indicated by different colors are initialized within a radius of $10^{-12}$ from the minimum in the direction of the Hessian eigenvector corresponding to the largest eigenvalue, for a 5-layer overparameterized scalar factorization of a random scalar. The value of $\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{w}^{*}))$ is computed using the closed-form expression derived in Theorem 21.

### v5

@@ D.2 Experiment Details and Additional Experiments
For the experiment, we first generate the layer dimensions randomly and then construct the optimal layers $\mathbf{W}_{1}^{*},\mathbf{W}_{2}^{*},\dots,\mathbf{W}_{L}^{*}]$ as Gaussian random matrices, with each entry sampled from $N(0,1)$ according to the generated dimensions. Then, we compute $\mathbf{M}$ or $m$ by $\prod_{j=1}^{L}\mathbf{W}^{*}_{j}$. We then perform the same experiments as in Figs. 14, varying the depth, dimensions, and initialization distance $r$ (as shown in Figs. 2-7). We note that oscillations occur along the eigenvector corresponding to the maximum eigenvalue of the Hessian. The dimensions of the factors, i.e., $d_{0},d_{1},\dots,d_{L}$}, in Fig. 6 are given by 1,9,4,8,24,16,17,11,21,3,22,3,3,15,3,18,17,16,5,12,1{1,9,4,8,24,16,17,11,21,3,22,3,3,15,3,18,17,16,5,12,1}, which implies $N=2421$, while the dimensions of the factors in Fig. 7 are given by 1,9,4,8,1{1,9,4,8,1}, which implies $N=293$.
(a) Normalized training error across iterations, i.e., $\mathcal{L}(\mathbf{w}_{k})/\lVert\mathbf{M}\rVert_{F}^{2}$.
(b) Normalized $\ell^{2}$ distance of $\mathbf{w}_{k}$ from the minimum $\mathbf{w}^{*}$, i.e., $\norm{\mathbf{w}_{k}-\mathbf{w}^{*}}_{2}^{2}/\norm{\mathbf{w}^{*}}_{2}^{2}$.
@@
(d) Normalized training error across iterations, i.e., $\mathcal{L}(\mathbf{w}_{k})/\lVert\mathbf{M}\rVert_{F}^{2}$.
(e) Normalized $\ell^{2}$ distance of $\mathbf{w}_{k}$ from the minimum $\mathbf{w}^{*}$, i.e., $\norm{\mathbf{w}_{k}-\mathbf{w}^{*}}_{2}^{2}/\norm{\mathbf{w}^{*}}_{2}^{2}$.
@@
Figure 3: GD dynamics with different step sizes indicated by different colors for the depth-$2$ matrix factorization of a random Gaussian matrix, where $\mathbf{L}\in\mathbb{R}^{10\times 30}$ and $\mathbf{R}\in\mathbb{R}^{20\times 30}$. The value of $\lambda_{\max}(\nabla^{2}\mathcal{L}(\mathbf{w}^{*}))$ is computed using the closed-form expression derived in Corollary 28.
(a) Normalized training error across iterations, i.e., $\mathcal{L}(\mathbf{w}_{k})/m^{2}$.
(b) Normalized $\ell^{2}$ distance of $\mathbf{w}_{k}$ from the minimum $\mathbf{w}^{*}$.
(c) Trajectories of GD on the contour map of the loss landscape.
(d) Normalized training error across iterations, i.e., $\mathcal{L}(\mathbf{w}_{k})/m^{2}$.
(e) Normalized $\ell^{2}$ distance of $\mathbf{w}_{k}$ from the minimum $\mathbf{w}^{*}$.
(f) Trajectories of GD on the contour map of the loss landscape.
Figure 4: GD dynamics with different step sizes indicated by different colors are initialized within a radius of $10^{-12}$ from the minimum in the direction of the Hessian eigenvector that corresponds to the maximum eigenvalue for a 15-layer overparameterized scalar factorization of a random scalar. The value of $\lambda_{\max}(\nabla^{2}\mathcal{L}(\mathbf{w}^{*}))$ is computed using the closed-form expression derived in Corollary 3.2.
(a) Normalized training error across iterations, i.e., $\mathcal{L}(\mathbf{w}_{k})/\lVert\mathbf{M}\rVert_{F}^{2}$.
(b) Normalized $\ell^{2}$ distance of $\mathbf{w}_{k}$ from the minimum $\mathbf{w}^{*}$, i.e., $\norm{\mathbf{w}_{k}-\mathbf{w}^{*}}_{2}^{2}/\norm{\mathbf{w}^{*}}_{2}^{2}$.
@@
(d) Normalized training error across iterations, i.e., $\mathcal{L}(\mathbf{w}_{k})/\lVert\mathbf{M}\rVert_{F}^{2}$.
(e) Normalized $\ell^{2}$ distance of $\mathbf{w}_{k}$ from the minimum $\mathbf{w}^{*}$, i.e., $\norm{\mathbf{w}_{k}-\mathbf{w}^{*}}_{2}^{2}/\norm{\mathbf{w}^{*}}_{2}^{2}$.
@@
Figure 5: GD dynamics with different step sizes indicated by different colors for general matrix factorization, $\mathbf{M}=\mathbf{L}\mathbf{R}^{\top}$, of a random Gaussian matrix, where $\mathbf{L}\in\mathbb{R}^{25\times 30}$ and $\mathbf{R}\in\mathbb{R}^{20\times 30}$. The value of $\lambda_{\max}(\nabla^{2}\mathcal{L}(\mathbf{w}^{*}))$ is computed using the closed-form expression derived in Corollary 28.
(a) Normalized training error across iterations, i.e., $\mathcal{L}(\mathbf{w}_{k})/m^{2}$.
(b) Normalized $\ell^{2}$ distance of $\mathbf{w}_{k}$ from the minimum $\mathbf{w}^{*}$, i.e., $\norm{\mathbf{w}_{k}-\mathbf{w}^{*}}_{2}^{2}/\norm{\mathbf{w}^{*}}_{2}^{2}$.
@@
(d) Normalized training error across iterations, i.e., $\mathcal{L}(\mathbf{w}_{k})/m^{2}$.
(e) Normalized $\ell^{2}$ distance of $\mathbf{w}_{k}$ from the minimum $\mathbf{w}^{*}$, i.e., $\norm{\mathbf{w}_{k}-\mathbf{w}^{*}}_{2}^{2}/\norm{\mathbf{w}^{*}}_{2}^{2}$.
@@
Figure 6: GD dynamics with different step sizes indicated by different colors are initialized within a radius of $10^{-9}$ from the minimum in the direction of the Hessian eigenvector that corresponds to the maximum eigenvalue for a 20-layer overparameterized scalar factorization of a random scalar. The value of $\lambda_{\max}(\nabla^{2}\mathcal{L}(\mathbf{w}^{*}))$ is computed using the closed-form expression derived in Corollary 3.2.
(a) Normalized training error across iterations, i.e., $\mathcal{L}(\mathbf{w}_{k})/m^{2}$.
(b) Normalized $\ell^{2}$ distance of $\mathbf{w}_{k}$ from the minimum $\mathbf{w}^{*}$, i.e., $\norm{\mathbf{w}_{k}-\mathbf{w}^{*}}_{2}^{2}/\norm{\mathbf{w}^{*}}_{2}^{2}$.
@@
(d) Normalized training error across iterations, i.e., $\mathcal{L}(\mathbf{w}_{k})/m^{2}$.
(e) Normalized $\ell^{2}$ distance of $\mathbf{w}_{k}$ from the minimum $\mathbf{w}^{*}$, i.e., $\norm{\mathbf{w}_{k}-\mathbf{w}^{*}}_{2}^{2}/\norm{\mathbf{w}^{*}}_{2}^{2}$.
@@
Figure 7: GD dynamics with different step sizes indicated by different colors are initialized within a radius of $10^{-12}$ from the minimum in the direction of the Hessian eigenvector that corresponds to the maximum eigenvalue for a 5-layer overparameterized scalar factorization of a random scalar. The value of $\lambda_{\max}(\nabla^{2}\mathcal{L}(\mathbf{w}^{*}))$ is computed using the closed-form expression derived in Corollary 3.2.